Date Time Explanation

09Aug09 16:25->18:05   There has been an unscheduled outage at 55 King St (iPrimus data centre) caused by equipment failure which impacted all our name lookup servers during the period shown. Issue is resolved.

27Jun09 21:15->23:30   There has been an unscheduled outage in the PowerTel network caused by equipment failure which impacted all our data feeds during the period shown. Issue is resolved.

06Aug07 07:30->08:15   There has been an unscheduled power outage at Global Centre in South Melbourne which impacted both our data feeds during the period shown. Issue is resolved.

23Jan07 10:30->30Jan07 10:00   aDSL service provider Comindico/SPTel has an issue with the main router in Melbourne causing users to be knocked offline and not being able to reconnect for a variable amount of time. It has been determined that the fault lies in the hardware and a replacement is being installed. Update: 30Jan2007 10:00am - Issue is resolved

15Nov06 16:00->16Nov06 11:30   aDSL service provider Comindico has introduced a new range of IP addresses which are allocated to end users. This new range has a routing issue which is preventing browsing to some sites (including ANZ, NAB, bureau of meteorology, techinfo, Bank of Queensland and others). 11:30am - Issue is resolved.

19Oct06 17:55->18:14   A routing issue at Prodigy (upstream) has caused slow browsing performance in the period shown.
All systems are normal again.

24Apr06 07:44->08:33   A power loss to the Briar Hill area has resulted in system shutdown (only) during the period shown.
Power to building fully restored at 11:28am

18Dec05 14:30->14:40   A short loss of external connectivity as we swapped core router to new unit as preparation for increased bandwidth.

03Sep05 11:10->19:00   An intermittent issue with the core router resulted in variable connectivity throughout the day as we have chased the problem. We believe it is now resolved and all services have returned to normal.

16Jul05 10:10->11:15   A hacker ran a 'ping flood' on the proxy machine rendering all external access to the net useless.

15Jul05 10:35->10:55   We have implemented a second fail over link to both increase bandwidth and provide some redundancy. The live test did not work and a routing loop resulted in no external connection for the period shown.

01Jul05 14:30->16:30   A Telstra data line fault has removed all external access to our servers. A temporary solution to restore connectivity is now in place until their technician arrives to replace the faulty equipment and restore the service.

27Apr05 20:30->28Apr05 10:00   Change to Telstra services knocked out all 56k indial lines between the period shown. All issues are fully resolved.

09Jan05 03:00->09:30   Browsing performance very slow due to a proxy server software issue - issue fully resolved, services returned to normal.

06Jan05 12:30->12:45   A short interruption to services whilst the main router at AAPT was replaced with a faster unit.

01Dec04 02:00->11:30   Mail filtering software upgrade caused outgoing mail not to be sent during this period. All backlog now cleared.

23Nov04 03:00->17:15   A customer's web page hosted on our server caused tremendous outgoing load resulting in a "Denial of Service" attack - all browsing slow during the period shown. Issue fully resolved.

14Oct04 21:15   Routing issue resolved on new router which has improved access speed.

07Oct04 07:30   New (more powerful) router installed in core network in preparation for bandwidth upgrade. Second machine to be installed at network border in next few weeks.

29Aug04 00:15->09:15   Main web server off line during the night whilst some upgrade work was performed.

16Aug04 18:30->17Aug04 10:30   There was a routing issue with one of our upstream providers during the period shown which prevented access to secure sites (banks, etc) and some other international sites. This issue has been resolved.

02Nov03 12:30->14:00   We restarted the proxy server several times in the period shown to implement a solution to the proxy server issue below. This problem should no longer be an issue with the network.

28Oct03 21:30->23:45   A problem with the proxy server software stopped browsing outside the network during the period shown.

27Oct03 16:00   A problem with duplicate IP addresses on an access server which was swapped in to the network as a replacement for a faulted unit led to 'dead' connections for some dial up customers - this has been resolved.

27Oct03 12:00   A problem with the 'online time usage web page' is being resolved - users cannot check their usage on line at this time.

25Oct03 13:00   There was a short duration outage of the RADIUS (login authentication system) server whilst we tested some changes to be made during the transfer the the new higher speed web server over the next few weeks.
We have also implemented some traffic shaping across our links to reduce the bandwidth available to spam e-mailers and make more bandwidth available for web browsing.

13Oct03 15:00   Have just commissioned a new more powerful proxy server replacing the previous unit. There was a five minute shutdown whilst the changeover occurred.

10Oct03 12:30->04:00   The performance of the web server was down as Chinese based spammers sent tens of thousands of mail messages to third parties. They and others have been blocked from doing so again.

06Oct03 11:30->07Oct03 09:05   All mail external to the techinfo network was rejected during the period shown as a black list we use to check for spam sources was faulting to reject all incoming mail. We no longer refer queries to this list.

06Oct03 11:30->11:35   All modems were shut down to allow installation of additional UPS and battery capacity. Several machine restarts were executed during the day as mail server and proxy server changes were updated.

29Sep03 18:30->18:40   The network was isolated from the internet as we replaced the gateway router with a new unit with more RAM installed to improve network performance.

22Sep03 17:30->18:30   The mail server machine was swapped for a complete new server today during the period shown. The service was stopped/started many times during the window causing some issues with mail delivery. All services returned to normal at the end of the period shown.

08Jun03 02:00->08:30   The proxy server faulted during the period shown resulting in no browsing access.

22May03 22:05   The level of spam filtering has been improved after an error in the filtering software configuration was discovered and rectified.

16May03 12:25->13:40   The mail server became overloaded and stopped receiving mail during the period shown after being attacked by several hundred thousand spam mail messages. The machine was cleaned out and restarted.

01Mar03 00:20->03:08   Total power failure in Briar Hill area during the period shown. Whole network running on UPS battery for first 1:5 hours, then forced system shutdown at 01:43am. Power restored at 02:07 and system restarted with a proxy server fault which was corrected and network became fully operational at 3:15am - network totally offline from 01:43am ->03:15am. All incoming mail will arrive.

21Feb03 11:32->12:45   Total power failure in Briar Hill area during the period shown. Whole network running on UPS battery for the duration. All services continued to function on battery power and no services were interrupted or impacted by the outage.

19Feb03 16:48->19:18   Total power failure in Briar Hill area during the period shown. Whole network running on UPS battery for the duration with supplemental power being supplied by long lead to an adjacent premises. All services continued to function on battery power and no services were interrupted or impacted by the outage.

24/25Jan03 22:30-07:30   A software problem with the proxy server stopped all browsing traffic and external access through the night. Problem has been solved and all services returned to normal.

24Oct02 16:30   We have changed the filtering software to another system and should be as aggressive in removing spam and less trouble to maintain.

04Oct02 18:00   We have implemented much more agressive spam and virus filtering software which analyses every incoming and outgoing mail message for typical spam content and virus attachments. These are deleted from the data stream. All users should receive very few spam messages as at this time.

18Sep02 17:40->20:39   Power failure in Briar Hill area during the period shown. Whole network running on UPS battery for the duration with supplemental power being supplied by very, very long lead to an adjacent premises.This is the longest power outage ever experienced. All services continued to function on battery power and no services were interrupted or impacted by the outage.

30Aug02 13:45   We have implemented a Realtime Blackhole List lookup service on the mail server to attempt to reduce the amount of SPAM (junk mail) being received by the server and passed on to our users.

26Aug02 12:40->15:30   Our upstream provider had an issue with routing outside their network to international links during the period shown. The issue is resolved.

17Aug02 20:28->21:20   Power failure caused network to operate on UPS during the period shown. No interruption to services. Power supply to one bank of 33k6 modems burnt out and replaced with spare. Calls to these 8 modems would have rang and not answered until power supply was replaced at about 10:00pm

07May02 20:25->22:50   The same 56k access machine shut down once again and would not accept calls.
Lines would have then been busy at various times for the period shown as the number of lines operational was halved.
The machine has been permanently replaced with an operational spare. This problem should not re-surface.

05May02 19:32->21:35   One 56k access machine shut down and would not accept calls.
Lines would have then been busy at various times for the period shown as the number of lines operational was halved.

07Feb02 18:10->18:15   The proxy server machine stopped briefly as the system hard drives were full of old log files.
The problem has been identified and has been remedied.

16Jan02 15:30->15:35   There was a 5 minute interruption to all services today whilst the router at AAPT was replaced with an upgraded model. All services have returned to normal.

06Aug01 05:00->06:00 will shut down mel-cor3 router to upgrade a card in the unit. This will stop all outgoing requests (mail, www, news, etc) for a 30 minute period during the window shown.

29Jul01 12:06->Mon 03:50   One of the new Southern Cross undersea fibres from Aus<-->USA has been cut about 20kms offshore from Sydney. This has reduced capacity to about 10% outgoing and incoming from USA.
As at 11:00pm, rerouting traffic has resulted in 50% capacity. Full capacity restored at 3:50am Monday.

20Jul01 00:01->10:30   We suffered a 'Denial of Service' attack resulting from the 'Code Red' virus on the main web server. Security patch has been installed and services returned to normal.

26Jun01 11:00->18:00   Outbound connectivity to internet has been slow for the afternoon due to a problem with's main DNS server being overloaded. This has resulted in mail being slow to exit our network to its final destination. Issue has been resolved.

04Apr01 18:55->19:45   Outbound connectivity to internet cut due to power supply failure in the main switch hub at's Melbourne network. Unit has been repaired by their technicians.

07Jan01 16:35->19:25   Connectivity to internet intermittent due to routing problem at Australian Netlink (another
Melbourne ISP) who incorrectly advertised a default route to the net out through their network.
Issue resolved.

02Jan01 08:30->12:00   Connectivity to internet disrupted due to routing problem at  Issue resolved.

18Dec00 16:00->   Routing problem with and Netspace has caused intermittent performance all afternoon and into the night until approx 21:30 when a fix was put in place. This problem has been with us since
Netspace changed upstream suppliers in early December. A permanent solution will be implemented
tomorrow afternoon after some consultation with engineers.

18Dec00 08:00-15:00   What happened to comsec/commsec this morning? Comsec have two domain names, & They changed internet hosting providers late on Sunday from UUNet
to Optus and sent through a redelegation request (a change to the 'internet phone book' saying our new
'phone number' is ''). They did this only for the name forgetting about This resulted in anyone trying to access them by the second name got an error message.
The second name has now also been changed over by CommSec this afternoon after they were shown the
error of their ways by other ISP's.

09Dec00 11:59   Server1 was accidentally taken offline by inexperienced operator on duty during the weekend at 11:59pm. Machine was returned to service at 08:09am 10Dec00.
Result: Network would not allow any one to login during the night.

29Nov00 07:35   All logins to network suspended for 5 minutes whilst database was reduced to a more manageable size.

20Nov00 17:30   Telstra has suffered a cut in one of their undersea cables to the Asia/Europe which impacts most of the internet in Australia. Response times to all US traffic will be slow until this is fixed.
Update 29Nov00: Expected date for cable repair is Wednesday 29Nov2000. Damage is 63km south of Singapore in a water depth of 25m. Cable runs from Singapore to Jakarta.

14Nov00 10:30 Additional battery pack installed into UPS extending total run time to 98minutes (estimated) up from 35minutes.

11Nov00 08:30-08:40 Shutdown 3 servers one at a time to swap over to new UPS system. All access servers (modems) will be swapped over on Monday morning (10 minute shutdown required then as well).

02Nov00 20:24-21:40 Power outage (apparently militant trade union action) forced a shutdown at 20:59
 - restarted network at 21:40. UPS battery backup system maintained power to network from

24Oct00 08:35 Telstra DDS Fastway service upgraded to 640kbps adding another 128kbps to the available bandwidth.

23Oct00 16:00-17:15 Power outage has forced a shutdown at 16:35 - restarted at 17:15.
 UPS battery backup maintained power to network for first 35minutes.

23Oct00 13:45-15:00 (our upstream provider) had routing issues with their main border router.

22Oct00 21:45-22:45 (our upstream provider) had routing issues with their main border router.
This resulted in intermittent operation for the period until their routers were reset.

07Oct00 15:30 Bloody hell, this day just keeps getting better!! - power supply in Bay Networks modem unit burnt out - replaced with idle VersaNet unit fitted with 24 modems (Bay has 60 modems).
Total 56k modem capacity down by 26 modems until Tuesday night when Bay should be fixed.
Update Tues. 10Oct00, 11:35am: Bay machine back online and all indial lines now available again.

07Oct00 13:30 Main Telstra DDS Fastway fibre optic link down - call logged, expect restoration late today.
There is no connectivity at all to "outside world" at the moment.
Update 5:05pm: Telstra are on their way to AAPT, Collins Street to replace faulty NTU unit.
Update 7:15pm: All services fully restored.

27Sep00 15:00 Reset Primary DNS server which had ceased to service DNS requests to 33k6 system for past
hour or so. Only effected those connecting on 33k6 rotary lines.

25Sep00 13:00 A power failure in the Briar Hill / Greensborough area has forced us onto battery backup power.
13:41 - Battery power now exhausted - system is shutting down
14:01 - Power restored to area by TXU - system restarting.
14:15 - all systems operational once again with the exception of one 56k machine (60 modems)
15:00 - second 56k machine operational once again after we carried out some maintenance.
  Network was offline for a total of 35 minutes from 13:41 to 14:15
  During this time if you dialled in, the following would have been the response:-
  33k6 system (9432-7222) - continuous ringing - no answer
  56k system (9433-4444) - busy signal

18Sep00 12:05 The routing problem which surfaced last week is with us again. (Access1/Ozemail hosted domains). A fix is in place and a permanent fix from Cisco is being sought and will be implemented tomorrow.
The problem is not in the Techinfo network but resides in the software in a router at Netspace. also have an routing issue with an upstream provider (Optus) which requires
  replacement hardware which will be installed in the early evening.

14Sep00 12:30 The routing problem to Access1/Ozemail hosted domains has been resolved (Comsec/Quicken/ANZwers have been inaccessible. Outgoing mail server overload problem has been resolved.

18Aug00 14:45 1 minute outage whilst new Cisco 2912XL 10/100Mbps switch was installed on network replacing D-Link unit..

09Aug00 07:50 Reduced number of records in online database for time usage. Records now start as at 01-May-2000

15Jul00 09:40 Restarted Telstra DDS Fastway link as NT1 unit had stopped functioning. Services returned to normal.

13Jul00 07:55 Rebooted Server1 twice after installation of Cold Fusion software for web site development.

12Jul00 17:00 Added 'expiry date' field to the account usage page so users can check their account expiry date.

11Jul00 07:40 Restarted one 56k machine after upgrade to latest V.90 code from supplier.
Both machines now at latest release, X17.0.2

12May00 07:45 Rebooted Server1 twice after upgrade to 256Mb RAM. All web services stopped temporarily.

23Apr00 11:45 Rebooted Primary DNS server as tape backup had faulted holding CPU utilisation at 100% for
past 4 hours. Those using the 33k6 system would have not been able to contact any other server either
internally eg. mail, news, etc or externally during this period.
The entire 56k system continued to function normally during this period.

09Apr00 08:40 Rebooted mail server several times over a 20 minute period whilst installing new video
card & drivers - apologies for interruption to service.

08/09Mar00 23:30 Power failure in the general Greensborough area has put us onto battery power.
System ran until 00:10 at which time we shut down with no power. Power restored at 00:50 ie. 40 minute shutdown of all services - this is becoming very, very tedious......

05Mar00 16:15 Power failure in the general Greensborough area has put us onto battery power.
System will run until 17:15 at which time we will be forced to shut down with no power.
Power restored at 17:30 ie. 15 minute shutdown of all services

29Feb00 06:00->9:00 (our upstream provider) had a problem with a border router causing
intermittent connectivity problems to all their services in Victoria. Problem was resolved around 9:00am.

03Feb00 15:15 State Government power restrictions forced a shutdown for 1 hour this afternoon.
Power was out from 2:15pm until 4:15pm. Network operated normally on power from UPS
system for first hour until batteries were drained.

17Jan00 14:30 Online time usage page error fixed. Change to 4 digit years on web server in late December for y2k compliance caused error in asp scripts which expect date to have 2 digit year, ie. '99' not '1999'

11Dec99 11:30->15:00 Web server / Radius server upgraded with larger hard drive.
Some interruption to services occurred during this period as machines were restarted many times.

26Nov99 21:25 Proxy server rebooted to check that problem has been resolved - apologies for the inconvenience.

26Nov99 18:30->19:15 Proxy server rebooted to remove 'buffer space' error but did not auto restart - problem will be looked into and solved later tonight.

04Nov99 10:45 The disconnection policy has been changed to give all users a maximum of 6 hours online in one session.  This replaces the previous 3 hour maximum session time.

02Nov99 12:20->13:40 were having routing issues within their network for the period specified.
  This resulted in very slow response to all www requests for the period.

16Oct99 11:30->14:20 AAPT exchange has had power failure knocking out our upstream provider We have no connectivity to the outside world during this period.

12Oct99 12:00->15:30 have routing issues with a border router. They have corrected the issue.

06Oct99 18:10 News server back online after correction of phone line / modem fault (has been offline for 3 days)

03Oct99 08:45->09:20 The RADIUS authentication database has been offline for the past 35 minutes whilst we
  compact the database back to a manageable size (250Mb is not considered manageable!).
  Nobody would have been able to login during this time (users would have had repeated requests
  for password). Apologies for the inconvenience.

02Oct99 07:10 We have used the last 10 minutes to carry out work on modifications to the RADIUS database
  having to take it offline.

08Sep99 11:00->13:00 Our upstream provider, '', have had problems with one of their routers causing extremely slow and intermittent response during the period shown. The problem is now reported by them to be fixed.

01Sep99 08:00 We have used the last 10 minutes to reduce the size of the online database. (now only May99->today)

15Aug99 19:09 Have fixed the FTP download page to correctly link to .exe files stored there.

15Aug99 18:00 Satellite connection to USA sites has been very slow all of today - has now been fixed by our upstream provider.

14Jul99 07:55 We have used the last 5 minutes to reduce the size of the online database. (now only March99->today)

23Jun99 08:55 We have used the last hour to reduce the size of the online database. No user logins during this time.

13Jun99 19:30 Web server software problem resolved and all domain hosted sites back online.

12Jun99 17:00 All servers have had hardware upgraded during the past 6 hours necessitating several shutdowns/restarts on each. Upgrades are completed although web server is not feeding Techinfo website currently. Will be fixed tomorrow.

03Jun99 08:00 Server1 shutdown for 10 minutes to replace noisy power supply fan. No logins from dial-in users allowed during this time.

28May99 16:40 Route filtering problem solved. Performance on www /games / IRC requests returned to normal.

28May99 09:10 Filtering from has slowed performance dramatically. Techs. working on problem.

26May99 20:00 Radius server issue resolved by installation of other (additional) software.

22May99 03:40 Radius server stopped authenticating all users connecting - server reset at 9:30am - all OK again.
Appears there is a fault which we will isolate.

20May99 03:39 Radius server stopped authenticating all users connecting - server reset at 8:30am - all OK again

07May99 01:24 Power failure caused fault in web server through night - reset at 07:30am.

30Mar99 18:30 Imported new active file for news server - around 13,500 newsgroups now available.

26Mar99 09:30 Radius server stopped authenticating all users connecting - machine reset - all OK again

24Mar99 10:30 Telstra DDS Fastway service being upgraded - 10 minute outage from now for line testing.

17Mar99 09:30 News server rebuilding history file - will take 12 hours and will be offline until late tonight.

10Mar99 08:00 Rebooted 56k machine to upgrade modem code to later version of V.90 - all users dropped off for 10 minutes.

25Feb99 14:45 Rebooted 56k machine - all users dropped off.

13Feb99 11:30 News server restarted and articles are flowing in again.

11Feb99 21:00 News server taken off line whilst file cache is deleted to start afresh. Index has become corrupt
and needs rebuilding.

27Jan99 13:00 Replaced motherboard in Proxy Server to stop random reboots of machine - no interruption to services.

27Jan99 08:30 Upgraded Bay Networks 56k machine to V.90 code - 3 minute interruption to 56k dial-in lines as machine restarted with new code. All 56k users should be able to achieve up to 44k connection speeds   (X2 modems as well as k56flex units after you upgrade your modem firmware to V.90 - refer to your  modem manual or for details on how to do this.)

23Jan99 15:00 Have completed the upgrade of the user database and migrated the mail server across to server2. All systems have returned to normal function. We will upgrade 56k system to V.90 next week.

09Jan99 07:05 APC UPS (battery backup) system placed back in service. 10 minute complete shutdown just completed.

07Jan99 17:00 APC UPS (battery backup) system running too hot. Advised by supplier to remove unit from operation. Entire network powered down for 3 minutes to carry out removal. New unit expected Friday 08Jan99. Another shutdown will be necessary then to replace unit. [AJC]

10Dec98 10:30 New (replacement) Bay Networks 56k box online - all dial-in services returned to normal (apologies to users for several restarts during the morning)   [AJC]

05Dec98 16:30 Alternate 56k system setup and operational. Will return to original system when warranty replacement unit arrives. 58 lines operational (2 down on standard). [AJC]

04Dec98 20:00 56k system has just blown a power supply - big smoke and lots of smell! - 56k service will probably be out until mid next week. Bay Networks have arranged to fly a new (replacement) unit in from USA over this weekend. [AJC]

03Dec98 15:45 Proxy server down from 14:15 -> 14:30pm today. CPU fan replaced and restarted. [AJC]

01Dec98 14:30 Our upstream provider had a problem with routing this morning in their Melbourne core system caused by failure of a device at Telstra. Degraded browsing performance experienced from 09:30 to 10:45. [AJC]